OpenStudy (anonymous):

For which system of equations would elimination be the simplest technique to choose?

OpenStudy (anonymous):

A. y = 2x and x + 3y = 5 B. 4x + 3y = 6 and 3y = 5 C. x + y = 7 and x - y = 12 D. 3x - 2y = 8 and 4x - 5y = 20

OpenStudy (anonymous):

A, simply replace y with 2x in the second equation and you can solve for x

OpenStudy (mertsj):

C Because if you just add the equations, y will be eliminated.

OpenStudy (mertsj):

@Lagrange That would be substitution.

OpenStudy (anonymous):

though it said substitution ....

OpenStudy (anonymous):

Thanks everyone! :)

OpenStudy (anonymous):

C. x + y = 7 and x - y = 12